A company is planning to purchase a machine that will cost $59,400 with a six-year life and no salvage value. The company expects to sell the machines output of 3,000 units evenly throughout each year. A projected income statement for each year of the assets life appears below. What is the payback period for this machine? Sales Costs: Manufacturing Depreciation on machine Selling and administrative expenses Income before taxes Income tax (30) Net income 141,000 $69,000 9,900 49,000(127,900) $ 13,100 (3,930) 9,170
Answer #1

Annual cash flow = 9170+9900 = 19070 per year

Payback period = Initial investment/Annual Cash flow

= 59400/19070

Payback period = 3.11 Years